REVISION DATE: 22-Feb-2018 10:08:37
In Line PSBS custom made by Mark, aligned on laser bench Pier 35, S/N 2. Test to eliminate PDBS and Fold mirror as source of instability Config10: 13_Feb-2018 to present. We took out the PDBS and the fold mirror. PD2 is connected to the through channel and PD1 has a 9V battery. FISH Fibers, SIRCUS Sphere, monitor with Si PD and SEI.
* 13-Feb-2018 22degC, 30% RH, no SEI today; chamber humidity went out of control and then recovered. * 14-Feb-2018 22degC, 30% RH; Labview crashed in the middle, creating data gap. * 15-Feb-2018 30degC, 30% RH; humidity control went out of spec at 19:38 then recovered; logging programs were finished by then. * 16-Feb-2018 20degC, 30% RH; (forgot I really wanted 22deg C). * 20-Feb-2018 20degC, 30% RH. * 21-Feb-2018 22degC, 40% RH.Figure 1: All data in this Configuration to date. PD signals are normalized to the monitor. Ratio PD1/PD2, Monitor, and normalized Signals all normalized to the initial value.
Figure 2: Zoomed in look at Day 1
Figure 3: Zoomed in look at Day 2
Figure 4: Zoomed in look at Day 3
Figure 5: Zoomed in look at Day 4
Figure 6: Zoomed in look at Day 5
Figure 7: Zoomed in look at Day 6